### v4

Abstract
Agentic large language model (LLM) training often involves multi-turn interaction trajectories that branch into multiple execution paths due to concurrent tool use, think-mode, sub-agent, context management and other runtime designs. As a result, the tokens produced by a single task naturally form a tree-structured token trajectory with shared prefixes, rather than a linear sequence. Existing training pipelines linearize such trajectories and treat each branch independently, leading to substantial redundant computation in both forward and backward passes. We derive that averaging the loss over all branches independently is algebraically identical to a per-token weighted loss, where each token’s weight equals the fraction of branches passing through it. The problem therefore reduces to computing the log-probability of every token in the prefix tree exactly once, with no repeated computation across shared prefixes: we propose DFS serialization of the tree, which visits every token exactly once, and adapt full-attention and SSM layers to ensure the resulting log-probabilities match independent per-branch calculation exactly. In practice, a single trajectory tree can be too large to fit in GPU memory; we therefore propose Tree Partitioning, a memory-efficient partitioning strategy that splits the tree into subtrees each fitting within GPU memory while preserving high prefix reuse. Together, these contributions form Tree Training, an efficient framework for training LLMs on tree-structured trajectories, achieving up to 6.2× end-to-end training speedup on dense and MOE models for both supervised fine-tuning and reinforcement learning.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Figure 1: Trajectory tree and DFS sequence. Left: a trajectory tree with $K=3$ root-to-leaf paths. Each node $n$ holds a token segment $T_{n}$; $g_{n}$ is the number of paths through $n$ (blue nodes are shared prefixes with $g_{n}>1$). Node text in red is model output (trained); black is user/environment input. Right: the DFS sequence places all nodes in depth-first order so that every token is computed exactly once, with no redundancy.
1 Introduction
Training Large Language Models (LLMs) in agentic scenarios involves multi-turn interactions with environments, where a single task frequently branches into multiple parallel paths. This branching is driven by tree-structured workflow designs (e.g., tree structured planning Yao et al. (2023); Zhang et al. (2024) and sub-agentsTeam et al. (2026)), concurrent tool invocations Kim et al. (2024); Abdelaziz et al. (2024), context management Xu et al. (2025); Packer et al. (2024); Zhong et al. (2023), as well as runtime behaviors such as think-mode (discarding/replacing intermediate reasoning tokens between turns, e.g. Figure 1), and retokenization drift Luo et al. (2025). These mechanisms imply that the context fed into turn t+1t{+}1 is often not exactly the concatenation of the tokens produced in previous turns (from turn $1$ to turn $t$); therefore, after finishing a task, the collected tokens cannot always be organized as a single list/tensor, but instead form a tree-structured token trajectory with shared prefixes across branches. As a result, training must process multiple sequences with substantial shared prefixes, leading to significant redundant forward/backward computation.
The decoder-only LLM’s generation is autoregressive and uses a causal attention mask. When we consider the inference scenario alone, prefix caching Kwon et al. (2023); Pope et al. (2022) can eliminate redundant prefix computation: the causal mask prevents suffix tokens from influencing prefix tokens, so identical prefixes produce identical key/value states, which can be reused to avoid recomputing the prefix in the forward pass. In training, the backward pass “transposes” the causal masking pattern: to compute the gradient of a prefix token, one must aggregate gradient contributions and intermediate activations from all suffix tokens under that prefix. Therefore, to avoid repeatedly recomputing prefix tokens, caching approaches would have to store suffix information for all branches, which drastically inflates GPU memory usage and exceed practical training budgets, making such approaches infeasible in practice.
To address the above challenges, we mathematically derive that the baseline loss—averaging over all branches independently—is algebraically identical to a per-token weighted loss where each token’s weight equals the fraction of branches passing through it. The problem therefore reduces to computing the log-probability of every token in the prefix tree exactly once, with no repeated computation across shared prefixes. We propose DFS serialization of the tree, which visits every token exactly once, and adapt full-attention and SSM layers to ensure the resulting log-probabilities match the baseline exactly.
We further redesign the training engine to natively ingest tree-structured data. We also address a key practical constraint: a single trajectory tree is often too large to fit entirely within GPU memory. We therefore propose Tree Partitioning, a memory-efficient partitioning strategy that splits a large tree into subtrees each fitting within memory while preserving a high prefix-reuse ratio.
Together, DFS serialization with model-layer adaptations, Tree Partitioning, and the training-engine redesign form Tree Training: an efficient framework for training LLMs on tree-structured trajectories for both SFT and RL, achieving up to 6.2× end-to-end training speedup in our experiments.
In summary, our contributions are threefold:
(1) We identify a pervasive yet overlooked property of agentic LLM training: trajectories naturally form overlapping tree-structured prefixes, creating substantial opportunities for computation reuse.
(2) We introduce Tree Training, a paradigm that executes each shared prefix once while achieving mathematically equivalent training to the baseline, via DFS serialization, targeted model-layer adaptations, Tree Partitioning, and the redesign of the training engine.
(3) We demonstrate consistent gains across dense and MOE DeepSeek-AI et al. (2024) models on diverse datasets, achieving average 6.2× end-to-end speedup—and in agentic RL scenarios—without compromising training fidelity or final model quality.
2 Related Work
Prefix caching is widely used to speed up autoregressive decoding by reusing identical prefixes in the forward pass Kwon et al. (2023); Pope et al. (2022); Liu et al. (2024); Cheng et al. (2024); Yao et al. (2025); Cheng et al. (2025); Qin et al. (2025). For training objectives where the shared part is purely a prompt, batching strategies such as prefix grouping can also reuse the prompt prefix computation without affecting gradients Liu et al. (2025); Wang and Hegde (2024). In addition, Goru et al. Goru et al. (2025) propose a custom-mask approach to deduplicate repeated reasoning tokens within a single linear conversation. However, it does not address prefix reuse across multiple trajectories, and its prefixes never aggregate loss from multiple branches. As a result, it is a special case of the agentic scenario where the backward pass remains standard.
Prior works have also recognized that agentic/RL generation often forms tree-structured rollouts, and exploit shared prefixes primarily to improve rollout efficiency, e.g., tree-search style rollouts and prefix reuse Hou et al. (2025); Li et al. (2025); Ji et al. (2025). However, the subsequent training stage is commonly executed on linearized samples: current training pipelines Luo et al. (2025); Zhang et al. (2025); Zhou et al. (2025) typically decompose a trajectory tree into separate linear segments and treat each segment independently, causing shared prefixes to be repeatedly recomputed in both forward and backward passes.
Sequence packing Krell et al. (2021) refers to concatenating multiple short sequences into a fixed-length training example to improve training efficiency. Our proposed tree partitioning method generalizes this idea: it not only concatenates sequences, but also introduces a strategy for efficiently segmenting overly long training data structured as a prefix tree (noting that a sequence is a special case of a prefix tree).
Recent hybrid LLM architectures interleave full-attention layers with State Space Model (SSM) layers. Gated Delta Net (GDN) Yang et al. (2025c) is a recent SSM variant that improves over earlier SSM architectures such as Mamba Gu and Dao (2024) and Mamba2 Dao and Gu (2024) by incorporating the delta rule, achieving better associative recall and in-context learning. GDN is adopted as the SSM component in hybrid models such as Qwen3.5 Team (2026). Unlike full-attention layers, SSM layers maintain a recurrent hidden state that is passed sequentially between tokens, which requires dedicated handling when processing tree-structured data.
Overall, unlike our work, these methods primarily target inference-time caching, prompt-only reuse, or single-conversation masking, and do not address the general case where a shared prefix must aggregate gradient contributions from multiple continuations.
Figure 2: SSM state routing. Sequential flow (left) feeds $n_{4}$’s state into $n_{2}$—wrong, since $n_{2}$ is a sibling of $n_{1}$, not a descendant of $n_{4}$. Tree routing (right) routes each chunk to its parent’s state: both $n_{3}$ and $n_{4}$ read $s_{n_{1}}$, while $n_{2}$ correctly reads $s_{n_{0}}$.
3 Tree Training
A trajectory tree is a rooted tree in which each node $n$ holds a token segment $T_{n}$. Each root-to-leaf path spells out a complete trajectory by concatenating the token segments of its nodes in order. Let $\mathcal{T}$ denote the set of all nodes and $\mathcal{P}=\{p_{1},\dots,p_{K}\}$ the set of $K$ root-to-leaf paths (one per leaf). An example with $K=3$ is shown in Figure 1. Goal. Process all tokens in the tree with each token computed exactly once (no redundant computation across shared prefixes) in a single forward and backward pass, and produce gradients that are mathematically equivalent to the baseline of running all $K$ paths independently and averaging their losses.
Derivation. The baseline sep-avg loss runs each of the $K$ paths independently:
$L_{\text{sep\_avg}}=\frac{1}{K}\sum_{k=1}^{K}\sum_{t\in p_{k}}\ell_{t}(\theta)$ (1)
where $\ell_{t}(\theta)$ is any per-token objective. For SFT, $\ell_{t}(\theta)=-\log p_{\theta}(y_{t}\mid x_{\leq t})$; for RL with policy gradient, $\ell_{t}(\theta)=-A_{t}\cdot\log p_{\theta}(y_{t}\mid x_{\leq t})$ where $A_{t}$ is the per-token advantage.
Eq. (1) iterates over paths then over tokens in each path. We swap the order: iterate over every unique token $t$ and count how many paths include it. Token $t$ belongs to node $n$; let $g_{n}=|\{p\in\mathcal{P}:n\in p\}|$ be the number of paths passing through $n$, and $g_{t}\triangleq g_{n}$ for $t\in T_{n}$. Token $t$ is therefore counted $g_{n}$ times in the double sum:
$\sum_{k=1}^{K}\sum_{t\in p_{k}}\ell_{t}(\theta)\;=\;\sum_{n\in\mathcal{T}}\sum_{t\in T_{n}}g_{n}\cdot\ell_{t}(\theta)\;=\;\sum_{t}g_{t}\cdot\ell_{t}(\theta)$ (2)
Substituting (2) into (1):
$L_{\text{sep\_avg}}=\sum_{t}\frac{g_{t}}{K}\cdot\ell_{t}(\theta)$ (3)
We therefore define the tree loss:
$\boxed{L_{\text{tree}}=\sum_{t}\frac{g_{t}}{K}\cdot\ell_{t}(\theta)}$ (4)
By linearity of differentiation:
$\frac{\partial L_{\text{tree}}}{\partial\theta}=\sum_{t}\frac{g_{t}}{K}\cdot\frac{\partial\ell_{t}}{\partial\theta}{=}\frac{\partial L_{\text{sep\_avg}}}{\partial\theta}\qquad$ (5)
This derivation is model-agnostic and objective-agnostic: since it relies only on the additive structure of $\ell_{t}$ across tokens and paths—not on its specific form—it applies to any per-token objective. The coefficient $g_{t}/K$ arises from equal weighting of $K$ paths ($w_{k}=1/K$); in the general case with path weights $\{w_{k}\}$, the per-token weight becomes $\sum_{k:\,t\in p_{k}}w_{k}$. For RL with policy gradient, $\ell_{t}(\theta)=-A_{t}\cdot\log p_{\theta}(y_{t}\mid x_{\leq t})$ with per-token advantage $A_{t}$; the $g_{t}/K$ factor remains the correct path-count normalization, while $A_{t}$ is already absorbed into $\ell_{t}$. Note that $L_{\text{tree}}$ and $L_{\text{sep\_avg}}$ are identical expressions—the substitution is purely algebraic; it does not yet impose any constraint on how $\ell_{t}(\theta)$ is computed in practice.
Problem reduction. The goal therefore reduces to: find a method that (1) computes each token exactly once with no redundant computation, and (2) ensures that $\ell_{t}(\theta)$ computed this way matches what it would be in a standalone per-branch forward pass—i.e., forward equivalence:
$\ell_{t}^{\text{tree}}(\theta)=\ell_{t}^{p}(\theta)\quad\forall\,p\in\mathrm{paths}(n_{t})$ (6)
We achieve this via DFS traversal of the token prefix tree, which serializes all nodes in depth-first order so each token appears exactly once, replacing the baseline serialization that repeats prefixes:
$X_{\text{base}}(i)=\operatorname{Concat}\bigl[\operatorname{token}(i),\,X_{\text{base}}(\operatorname{child}(0,i)),\\ \operatorname{token}(i),\,X_{\text{base}}(\operatorname{child}(1,i)),\;\dots\bigr]$ (7)
with:
$X_{\text{DFS}}(i)=\operatorname{Concat}\bigl[\operatorname{token}(i),\,X_{\text{DFS}}(\operatorname{child}(0,i)),\\ X_{\text{DFS}}(\operatorname{child}(1,i)),\;\dots\bigr]$ (8)
How to enforce (6) on the DFS-serialized sequence—i.e., ensuring each token’s log-probability matches its per-branch value despite the reordering introduced by DFS traversal—is described in Section 3.1.
3.1 Implementation
Loss Weights.
For each token $t$ in the DFS-serialized sequence, its per-token loss is weighted by $g_{t}/K$ before reduction, where $g_{t}$ is the number of root-to-leaf paths passing through token $t$’s node and $K$ is the total number of leaves. This is applied as a single element-wise multiplication on the per-token loss tensor—requiring no modification to the backward pass and negligible overhead.
With the following two adaptations, the forward equivalence condition (6) is satisfied and the gradient equivalence holds.
Attention Mask.
We construct a tree attention mask such that token $i$ attends to token $j$ iff (1) $j\leq i$ and (2) the nodes of $i$ and $j$ lie on a common root-to-leaf path. This prevents cross-branch attention. We implement a GPU kernel based on Flash Attention V3 Shah et al. (2024); Wang et al. (2025) with node-level shared-prefix masking. The attention pattern is illustrated in Figure 3.
Figure 3: Tree attention mask. Left: the $6\times 6$ attention matrix for the DFS sequence; a cell is filled iff the query token can attend to the key token. Tokens from different branches (e.g. $n_{3}$ and $n_{2}$) are mutually blocked. Right: the corresponding tree structure; the dashed red arrow shows a blocked cross-branch path.
Position Embedding.
Each token must receive the same position ID as it would in its standalone per-path sequence. A token at offset $j$ within node $n$ is assigned:
$\mathrm{pos}(t)=\sum_{n^{\prime}\in\mathrm{anc}(n)\setminus\{n\}}|T_{n^{\prime}}|+j$ (9)
where $\mathrm{anc}(n)$ denotes the ancestors of $n$. Sibling nodes at the same depth share the same position range, matching the per-path baseline and ensuring RoPE Su et al. (2023) produces identical results.
SSM Layers (GDN).
For hybrid models that interleave full-attention with Gated Delta Net (GDN) Yang et al. (2025c) SSM layers, the attention mask alone is insufficient: GDN layers maintain a recurrent hidden state passed sequentially between chunk boundaries, so a naïve DFS layout would carry the state from one branch into the next sibling branch. We apply two fixes. We treat each tree node’s token segment as one chunk—the unit of SSM state transfer. (i) State routing: each chunk $c$ reads its initial SSM state from its parent chunk $\pi(c)$ rather than the preceding chunk in DFS order:
$\mathbf{h}_{c}^{(0)}=\mathbf{h}_{\pi(c)}^{(\mathrm{end})},\qquad\mathbf{h}_{-1}^{(\mathrm{end})}\triangleq\mathbf{0}$ (10)
DFS order guarantees $\pi(c)<c$, so parent states are always ready. Sibling chunks share the same parent state tensor, so autograd accumulates their gradients there automatically. Figure 2 illustrates the difference between sequential and tree state routing.
(ii) Causal convolution: SSM layers include a causal conv1d with kernel size $K_{\mathrm{conv}}$, which requires the preceding $K_{\mathrm{conv}}{-}1$ tokens as left context. In the DFS sequence, the tokens immediately before a child chunk belong to a sibling branch rather than the ancestor path, so using them as context is incorrect. Instead, each chunk $c$ uses $\mathbf{ctx}_{\pi(c)}$—the last $K_{\mathrm{conv}}{-}1$ output tokens of its parent chunk $\pi(c)$, saved after processing the parent—as its left context:
$\mathbf{y}_{c}=\mathrm{Conv}\!\left([\,\mathbf{ctx}_{\pi(c)}\,;\,\mathbf{x}_{c}\,]\right)_{[K_{\mathrm{conv}}-1:]}$ (11)
This replicates the context each token would see in an independent per-branch forward pass. Figure 4 illustrates this for $K_{\mathrm{conv}}=3$.
Figure 4: SSM causal convolution ($K_{\mathrm{conv}}=3$). When processing $n_{2}$ (child of $n_{0}$), the correct conv window is formed by prepending conv_state[$n_{0}$]—the stored last $K_{\mathrm{conv}}{-}1$ effective tokens of the parent—as left context, bypassing the DFS-adjacent sibling tokens ($n_{4}$) that occupy the preceding positions.
3.2 Tree Partitioning
Figure 5: Tree Partitioning under memory constraints. When the full tree (83k tokens) exceeds GPU memory (60k), our method partitions it into two sub-sequences (102k total) vs. the baseline’s 164k.
In practice, fitting an entire trajectory tree into GPU memory is often infeasible; therefore, we design an efficient packing algorithm that partitions a large computation tree into a sequence of memory-feasible subtrees. The goal is to keep each subtree within a memory limit $C$ while maximizing shared-prefix reuse, thereby minimizing the overall training cost.
Theoretically, the optimal partitioning requires dynamic programming coupled with bin packing, which is NP-hard and computationally prohibitive for large-scale trees due to an exponential state space. Implementation details are provided in the appendix A.
To balance efficiency and performance, we instead adopt a greedy heuristic DFS algorithm that approximates the optimal packing while scaling linearly with tree size. This approach follows three guiding principles: it prioritizes allocating the deepest leaves first (as they contribute most to the total trajectory length), groups leaves of similar depths within each subtree to improve packing homogeneity, and traverses the tree in depth-first order, initiating a new traversal whenever the accumulated length exceeds capacity $C$.
To demonstrate the efficiency of this approach, we present a concrete example in Figure 5. Consider a tree-structured dataset containing 4 trajectories with a total of 83k tokens. Assuming a GPU memory limit of $C=60\text{k}$, standard flattening would yield 4 separate sequences totaling 164k tokens due to redundant prefixes. In contrast, our algorithm partitions the tree into two sequences (fitting within $C$), resulting in a total of only 102k tokens. This significantly reduces both memory usage and computational overhead compared to the baseline.
Figure 6: Real agentic trajectory trees and their overlap characteristics. The upper row shows representative trees from multi-turn agentic RL rollouts with Low, Medium, and High Overlap. The lower row plots active trajectory counts over length; the area ratio reflects the theoretical token reuse ratio.
Figure 7: End-to-end training speedup and loss comparison using real world rollout data with think mode turned on. Results for the MoE model Qwen3-30B Yang et al. (2025a) are shown on the left, and for the dense model Qwen3-32B on the right. For each model, the top plot illustrates the training speedup achieved by Tree Training, while the bottom plot displays the average relative error in loss.
3.3 Does Tree Training introduce bias in the gradient updates?
Tree Training does not introduce bias during training. This is because our method does not construct trees across different global batches. During training, each global batch is a self-contained tree. The data shuffle operation is applied only between these complete tree samples and never disrupts the internal structure within a tree (as illustrated in Figure 1, where a Global Batch Size (GBS) of 3 corresponds to three independent tree structures).
The fundamental reason is: The set of trajectories forming a single tree does not come from multiple rollouts; instead, the prefix-tree structure is naturally generated from a single rollout. Consequently, the tree partitioning is organized and processed entirely within one gradient accumulation step. It never involves packing trajectories from different global batches.
4 Experiments
We evaluate Tree Training along three dimensions: gradient correctness, efficiency under realistic memory constraints, and end-to-end training performance. Across all settings, reusing shared prefixes from overlapping trajectories substantially reduces redundant computation while preserving equivalence to standard sequence-based baselines.
4.1 Metrics
To clearly separate dataset structure from runtime effects, we define two complementary metrics.
Potential Overlap Ratio (POR) quantifies the theoretical upper bound of computation reuse by measuring the redundancy inherent in shared prefixes. We define POR as the ratio of overlapping tokens to the total token count of the flattened dataset:
$\text{POR}=1-\frac{N_{\text{tree}}}{N_{X_{base}(root)}}$ (12)
Here, $N_{\text{tree}}$ denotes the token count in the prefix tree, while $N_{X_{base}(root)}$ represents the token counts of the serialized tree produced by the baseline method described in Equation (1). Using the example from Figure 5, the POR is calculated as $1-\frac{83\text{k}}{164\text{k}}\approx\mathbf{49.4\%}$, indicating that nearly half of the baseline computation is redundant.
Effective Reuse Ratio (ERR) evaluates the actual efficiency of computation reuse under memory constraints. It is defined as the fraction of redundant tokens actually eliminated by the optimization:
$\text{ERR}=1-\frac{N_{\text{pack}}}{N_{X_{base}(root)}}$ (13)
Here, $N_{\text{pack}}$ represents the total token count after partitioning the tree. For the scenario in Figure 5, the ERR is calculated as $1-\frac{102\text{k}}{164\text{k}}\approx\mathbf{37.8\%}$, demonstrating significant savings despite the memory constraint.
4.2 Evaluation Setups
Figure 8: End-to-end training speedup of Tree Training across datasets with varying POR. Each subfigure reports the relative reduction in total training time of Tree Training compared to the baseline. (a) synthetic datasets where the full tree fits in GPU memory, (b) synthetic datasets requiring tree partitioning under memory constraints.
We evaluate the end-to-end training efficiency by comparing the total time required to process equivalent data between Tree Training and the baseline described in Section 3. This baseline aligns with the current standard practice for training on branched trajectories, as most existing training frameworks only support taking batch of list (tensor) as training inputs.
All experiments were conducted on clusters of 64 NVIDIA Hopper GPUs using Megatron-Core Shoeybi et al. (2020) for distributed training. Sequence Packing Krell et al. (2021) is used in our baseline setting, representing the standard batching strategy used in large-scale LLM fine-tuning and RL training.
4.3 Numerical Precision
While the derivation establishes exact mathematical equivalence, hardware floating-point arithmetic introduces unavoidable small numerical deviations in practice. The DFS-serialized forward pass accumulates intermediate results in a different order than independent per-branch passes, and cuBLAS may select different GEMM kernels for different tensor shapes, both causing non-associativity errors inherent to bfloat16 arithmetic. These are not implementation bugs but fundamental hardware-level effects, and as shown in the next section, they do not affect training dynamics in practice.
4.4 Speedup and Correctness on real-world agentic trajectories
To assess real-world applicability, we collect agentic RL rollouts from multi-turn reinforcement learning tasks. Figure 6 presents representative realistic agentic trajectory trees (top) alongside their corresponding token-overlap distributions (bottom). These trajectories are collected from software engineering tasks (SWE-smith Yang et al. (2025b)), executed via agentic scaffolds Claude Code 4. Most branching behaviors of the left two trees are caused by concurrent tool execution and retokenization drift Luo et al. (2025), while the right tree is caused by the think mode. These examples illustrate a broad spectrum of prefix sharing, ranging from low to high POR (from 28.0% to 88.7%), and reveal that real agentic datasets tend to exhibit sparse, unbalanced trees with highly variable reuse across different depths and branching factors.
To validate the correctness and speedup of our method, we illustrate the loss comparison and step-time comparison on real-world agentic trajectories as shown in Figure 7:
(1) Speedup (Top Panel): The theoretical speedup upper bound $\frac{1}{(1-\text{POR})}$ for this dataset is 6.5×. We achieved average realized speedups of 6.3× (32B Dense) and 6.2× (30B MoE). This demonstrates that our system incurs negligible overhead, capturing over 95% of the theoretical potential. The step-wise fluctuation (2×–10×) faithfully reflects the varying intrinsic POR of dynamic tree samples.
(2) Correctness (Bottom Panel): Tree Training reproduces the baseline’s loss and gradient behaviors, with numerical deviations consistently maintained within acceptable precision tolerance. The training loss curves for Baseline (Yellow) and Ours (Green) overlap perfectly. The deviation (Red dashed) is negligible (¡1%), confirming strict mathematical equivalence.
4.5 Speedup of datasets with varying POR
For controlled experiments, we construct synthetic datasets with POR values from 20% to 92%, while keeping the number of leaf trajectories and total tokens constant across settings. This setup allows us to systematically study how the overlap degree influences computation reuse and throughput.
As shown in Figure 8, Tree Training consistently reduces total training time across all conditions, with the improvement magnitude increasing monotonically with the Potential Overlap Ratio (POR). In the ideal setting where full trees fit in memory, Tree Training achieves up to 8.7× speedup by eliminating redundant computation through shared prefix reuse.
4.6 Memory footprint
In our implementation, even when large trees must be split (achieved via the proposed Tree Partitioning method), our performance remains close to $\frac{1}{(1-\text{ERR})}$, demonstrating that our system incurs negligible runtime overhead. The additional tensors required by Tree Training total only 1.2 MB on Qwen3-32B Yang et al. (2025a), compared to the 64000 MB minimum activation memory, confirming negligible memory overhead.
4.7 The gain of training on all tokens of trajectory trees
Training on all tokens of trajectory trees, compared to training on only a single trajectory (in our baseline experiment, we select the longest trajectory as common practice), leads to a clear performance gain, especially for thinking models.
On Terminal Bench 2.0Merrill et al. (2026), we use a cold-started Qwen3-32B Yang et al. (2025a) as the base checkpoint for RL training. The model trained on the full tree (with think-mode turned on) achieves a score (avg@4) of 28.8, whereas the baseline score is 20.9.
5 Conclusion
We propose Tree Training, a performance optimization framework during the training phase. It aims to minimize redundant computation and boost efficiency while strictly maintaining the gradient values of the parameters undergoing optimization. In contrast to existing prefix reuse schemes, our Tree Training is not only suitable when the prefix is just a prompt that does not contribute to gradients, but also when the prefix is involved in gradient updates.
Appendix A Tree Partitioning DP solution
Instead of repeatedly processing shared prefixes, we merge trajectories into a tree structure that explicitly captures shared computation. However, fitting the entire trajectory tree into GPU memory is often infeasible, motivating the need for an efficient partitioning strategy. Tree Partitioning formulates this as optimizing how to split the computation tree into subtrees under a total trajectory-length budget $C$, ensuring that each batch stays within capacity while maximizing reuse of shared prefixes to minimize overall training cost.
A.1 Single-Path Packing
Within the trajectory tree, each node represents a token segments shared across multiple rollouts, such as a common prompt or an intermediate environment feedback. Selecting an internal node $u$ and using its root-to-$u$ path as a shared segment allows all descendant leaves to reuse this computation. We first analyze a simplified setting where each training step establishes only a single shared path within the trajectory tree. In this case, a training step selects one node $u$ whose shared and residual trajectory lengths jointly fit within the global capacity $C$. Dynamic programming is employed to enforce this feasibility constraint and compute the optimal selection efficiently.
Notation and constraint
Let $T=(V,E)$ denote the trajectory tree with root $r$. Each edge $e=(x,y)\in E$ has a length $l(y)$ corresponding to the token length of the child sub-trajectory $y$. For the root $r$, $l(r)$ denotes the segment length of the root itself. For nodes $p,q\in V$, define $d(p,q)$ as the total token length along the path from $p$ to $q$. For an internal node $u\in V$, we define:
  • $L(u)=d(r,u)$ is the length of the shared prefix $r\to u]$.
  • $\mathcal{L}(u)$ is the set of leaves in the subtree rooted at $u$, and $n_{u}=|\mathcal{L}(u)|$ its cardinality.
  • $R(u)=\sum_{\ell\in\mathcal{L}(u)}d(u,\ell)$ is the total residual length under $u$.
A node $u$ is feasible if it satisfies
$L(u)+R(u)\leq C.$ (14)
The length saving achieved by selecting $u$ as a shared node is
$\mathrm{S}(u)=(n_{u}-1)\,L(u).$ (15)
If $u$ is infeasible, it cannot be chosen as a shared node in any valid training step.
DP definition with feasibility
Define $DP(u)$ as the maximum total length saving achievable while covering the entire subtree rooted at $u$, using only feasible shared-node selections. If it is impossible to cover all leaves within $u$’s subtree under the constraint, we set $DP(u)=$-$\infty$.
The recurrence is:
$DP(u)=\begin{cases}0,&\text{if $u$ is a leaf},\\[2.0pt] \max\Bigg\{\mathbf{1}_{\text{f}(u)}\cdot(n_{u}-1)L(u),\\ \sum_{v\in\text{child}(u)}DP(v)\Bigg\},&\text{otherwise}.\end{cases}$ (16)
Here $\mathbf{1}_{\text{f}(u)}=1$ if $u$ is feasible and 0 otherwise. The term $\sum_{v\in\text{child}(u)}DP(v)$ is -$\infty$ if any child $v$ has $DP(v)=$-$\infty$. Quantities $L(u)$, $n_{u}$, and $R(u)$ can be precomputed in a single post-order DFS traversal in $O(|V|)$ time.
Figure 9: Comparison between single-path and multi-path tree partitioning. Step 1 packs the shared prefix $r\!\rightarrow\!u\!\rightarrow\!v_{1}$, and Step 2 packs $r\!\rightarrow\!u\!\rightarrow\!v_{5}$ separately. The optimal strategy instead treats $r\!\rightarrow\!u\!\rightarrow\!\{v_{1},v_{5}\}$ as a hierarchical shared prefix, enabling greater computation reuse, as discussed in section A.2.
Intuitively, the dynamic program either places a shared prefix at node $u$, or recursively delegates coverage to its child subtrees. However, restricting each training step to a single shared prefix may fail to achieve maximal reuse, especially when the capacity $C$ is relatively large. As illustrated in Figure 9, packing $r\!\rightarrow\!u\!\rightarrow\!v_{1}$ and $r\!\rightarrow\!u\!\rightarrow\!v_{5}$ in separate steps leads to redundant computation, whereas the optimal multi-path strategy jointly reuses $r\!\rightarrow\!u\!\rightarrow\!\{v_{1},v_{5}\}$ as a hierarchical shared prefix. We therefore extend the formulation to the multi-path setting, where multiple shared prefixes can be activated simultaneously within a single training step.
A.2 Multi-Path Tree Partitioning
We now generalize Tree Partitioning to the multi-path setting. Selecting an internal node $u$ activates multiple shared paths: the root-to-$u$ path becomes the upper-level shared trajectory, and for each child of $u$, additional root-to-child trajectories may be nested and reused. In this way, shared sub-trajectories can branch and overlap, covering multiple subtrees in one step and further reducing the total length, as some sub-trajectories are shared across more leaves.
This leads to a dynamic programming formulation in which each subtree yields a set of candidate states. Parent states are constructed via two operations: (Lift) propagating child states upward by adding the edge length, and (Bin packing) combining the lifted demands into feasible traversals under capacity $C$.
Notation
We reuse the notation from the single-path setting. The local remaining capacity at node $u$ is
$C_{u}=C-L(u).$ (17)
State
For each node $u$, we maintain a candidate set
$\mathcal{A}_{u}=\big\{(\mathbf{b},\,\mathrm{cost}(\mathbf{b}))\big\},$ (18)
where $\mathbf{b}=(b_{1},\dots,b_{k})$ is a non-decreasing vector with $0\leq b_{i}\leq C_{u}$. Here, $k$ is the number of traversals that arrive at $u$, each $b_{i}$ represents the capacity already used in that traversal upon entering $u$, and $\mathrm{cost}(\mathbf{b})$ denotes the minimum cumulative length required to cover all leaves in $\mathcal{L}(u)$ under this configuration.
Initialization
If $u$ is a leaf, then
$\mathcal{A}_{u}=\{\,((0),\,0)\,\}.$ (19)
Recurrence (child $\to$ parent)
Suppose $u$ has children $\{v_{1},\dots,v_{m}\}$ with corresponding candidate sets $\mathcal{A}_{v_{j}}$. We construct $\mathcal{A}_{u}$ in two steps:
  • Lift: For each child $v$ and each entry $(\mathbf{t},\,\mathrm{cost}_{v}(\mathbf{t}))\in\mathcal{A}_{v}$, define
    $\mathbf{t}^{(u)}=\mathbf{t}+\operatorname{w}(v)\cdot\mathbf{1},$ (20)
    and keep only those with $\max(\mathbf{t}^{(u)})\leq C_{u}$. Each component of $\mathbf{t}^{(u)}$ is treated as an item and inserted into a multiset $S$, with its associated cost contribution $\mathrm{cost}_{v}(\mathbf{t})$.
  • Bin packing: For each feasible packing of $S$ into bins of capacity $C_{u}$, let the bin loads be $b_{1},\dots,b_{k}$ (sorted non-decreasingly). This defines the parent state $\mathbf{b}=(b_{1},\dots,b_{k})$. For each child $v$, let $t_{v}$ be the number of bins containing items lifted from $v$. The total cost is
    $\mathrm{cost}(\mathbf{b})\;=\;\sum_{v}\mathrm{cost}_{v}(\mathbf{t}_{v})\;+\;\sum_{v}\operatorname{w}(v)\cdot t_{v}.$ (21)
    Insert $(\mathbf{b},\,\mathrm{cost}(\mathbf{b}))$ into $\mathcal{A}_{u}$, keeping only the minimal cost for identical $\mathbf{b}$.
Termination
At the root $r$, we obtain $\mathcal{A}_{r}$. The optimal total length is
$\mathrm{Cost}=\min_{(\mathbf{b},c)\in\mathcal{A}_{r}}c.$ (22)
Complexity and Approximation
The multi-path dynamic programming formulation provides a theoretically optimal solution but is computationally expensive in practice. The state space may grow exponentially in the worst case, since each child contributes multiple candidate vectors and their combinations grow multiplicatively across children. Furthermore, the bin-packing step on the multiset $S$ is locally NP-hard.
Thus, the exact DP is tractable only for small to medium trees but remains valuable as a ground truth for evaluating approximations. For larger instances, we complement the exact DP with a heuristic algorithm that approximates the bin-packing step while scaling efficiently to large trees. In practice, the heuristic algorithm follows three guiding principles: it prioritizes allocating the deepest leaves first (as they contribute most to the total trajectory length), groups leaves of similar depths within each subtree to improve packing homogeneity, and traverses the tree in depth-first order, initiating a new traversal whenever the accumulated length exceeds capacity $C$.